OpenStudy (megannicole51):

does the series converge or diverge? (2n)!/(n+3)! n=1 to infinity

hartnn (hartnn):

did you try ratio test ?

OpenStudy (megannicole51):

let me do that now:)

OpenStudy (megannicole51):

An=(2n)!/(n+3)! An+1=(2n+1)!/((n+3)+1)! ??

hartnn (hartnn):

correct

OpenStudy (megannicole51):

((2n+1)!/((n+3)+1)!)/((2n)!/(n+3)!)

hartnn (hartnn):

wait, 2n -------> 2 (n+1) = 2n+2

OpenStudy (megannicole51):

oh okay! i thought something would be wrong haha

hartnn (hartnn):

and use the fact that (x+1)! = (x+1)x!

hartnn (hartnn):

(2n+2)! = (2n+2)(2n+1)2n! (n+4)! = (n+4)(n+3)!

OpenStudy (megannicole51):

((2n+2)!/(2n)!)*((n+3)!/(n+3)+1)!) how do i simplify this?

OpenStudy (megannicole51):

wait what?

hartnn (hartnn):

did you know this ? (x+1)! = (x+1)x! ?

OpenStudy (megannicole51):

nope:/

hartnn (hartnn):

how will you get say umm, 4! =.. ?

OpenStudy (megannicole51):

4*3*2*1

hartnn (hartnn):

3! = 3*2*1 so, 4! = 4*3! this is true in general

hartnn (hartnn):

n! = n (n-1)!

OpenStudy (megannicole51):

makes sense:)

hartnn (hartnn):

so, (2n+2)! = (2n+2) (2n+1)!

hartnn (hartnn):

and (2n+1)! = (2n+1)2n!

hartnn (hartnn):

so you got this ? (2n+2)! = (2n+2)(2n+1)2n! (n+4)! = (n+4)(n+3)!

OpenStudy (megannicole51):

i think so.....also, is (n+3)+1) simplified to just n+4?

hartnn (hartnn):

yes, thats just (n+1)+3 = n+4

OpenStudy (megannicole51):

ok just checking

OpenStudy (megannicole51):

do i have to use that rule for (2n)! and (n+3)! too?

hartnn (hartnn):

those will get cancelled , right ?

OpenStudy (megannicole51):

yes...i see it now lol

OpenStudy (megannicole51):

(2n+2)(2n+1)/n(n+4)

hartnn (hartnn):

how does "n" come in the denominator ?

OpenStudy (megannicole51):

typo sorry!

hartnn (hartnn):

no problem can you evaluate lim n->infinity (2n+2)(2n+1)/(n+4)

OpenStudy (megannicole51):

yeah hold on

hartnn (hartnn):

*holding tight*

OpenStudy (megannicole51):

haha....had to get my drink! isnt it infinity bc as n in the denominator gets bigger the entire den. gets bigger? approaching infinity?

OpenStudy (megannicole51):

so it diverges?

hartnn (hartnn):

absolutely correct!

hartnn (hartnn):

yes

OpenStudy (megannicole51):

yay!!!! omgsh thank you so much! i'm so going to get this problem right if its on my exam!

hartnn (hartnn):

glad to hear! ok, just an algebraic way of solving the limit multiply numerator and denominator by 1/n ^2 as n-> infinity , 1/n -> 0, 1/n^2-> 0 so you will get a 0 in the denominator , implies lim = infinity

OpenStudy (megannicole51):

okay!! now i have a question....how do u know when to n(n-1)!? bc we didnt use it on 2n! and (n+3)!

hartnn (hartnn):

we had (2n+2)! in numerator and 2n! in denominator, right ? its like having 5! in numerator and 3! in denominator, so the only thing we can use is n! = n(n-1)! to simplify

OpenStudy (megannicole51):

okay so u just look at it and think well if i simplify this by using the formula it will cancel that out!

hartnn (hartnn):

correct ... and when you do few more problems like this, you'll get what to do automatically in other similar problems :)

OpenStudy (megannicole51):

can u work on another with me? i already worked it out but idk if its correct

OpenStudy (megannicole51):

im studying for my exam this wednesday and there isnt an answer key to my practice problems so idk if im getting any of these correct or not:(